OpenStudy (gigi1103):

Which shows three ordered pairs that are all solutions of the equation y = x + 4? A. (1, 4), (2, 6), and (3, 7) B. (1, 4), (2, 8), and (3, 12) C. (3, 7), (5, 9), and (7, 10) D. (3, 7), (5, 9), and (7, 11)

OpenStudy (gigi1103):

@tomorrow

OpenStudy (gigi1103):

@MakaylaTracy

OpenStudy (gigi1103):

@ElonaSushchik

OpenStudy (gigi1103):

can you help

OpenStudy (anonymous):

each y must be 4 more than it's x in the ordered pairs given. check some and see which fit.

OpenStudy (gigi1103):

i don't get it

OpenStudy (alexandervonhumboldt2):

substitute (x, yI values nto the equations and check

OpenStudy (anonymous):

put each ordered pair from the solutions in your equation and see which ones are correct.

OpenStudy (gigi1103):

@Here_to_Help15

OpenStudy (anonymous):

For example: y= x +4 replace y and x with the an ordered pair such as (1, 4) which will be 4= 1+4.... 4 = 5 thats wrong...now you get it?/

OpenStudy (gigi1103):

is it a

OpenStudy (gigi1103):

yes i get it

OpenStudy (anonymous):

nice job!!!!

OpenStudy (gigi1103):

but none make four!!!!

OpenStudy (gigi1103):

ummm is it a

OpenStudy (gigi1103):

wait no its d i get it thaanks you guys

OpenStudy (anonymous):

Yes D. is correct :)

OpenStudy (gigi1103):

one more question

OpenStudy (anonymous):

okay